It took a total of 10 hours to hike from the base of a mountain to the top and
back by the same route. Going up speed was 2 miles per hour. Going down speed was 3 miles per hour. How far from base to top?
A) 6 B) 10 C) 12 D) 14 E) 24

Answer by palanisamy(496)

Let the distance from the base to the mountain top = x
Going up speed was 2 miles per hour
Time taken = x/2
Going down speed was 3 miles per hour.
Time taken = x/3
Total time taken is x/2+x/3 = 10
(3x+2x)/6 = 10
5x = 60
x = 60/5
x = 12 miles